1) A 178 g copper calorimeter contains 601 g of water at 19.2°C. Steam at 100°C is bubbled through the water. After a while, the mass of water in the calorimeter has increased by 9.5 g, and the temperature is 30.3°C.

Use this information to determine an approximate value for the latent heat of vaporization of water( Mg/kg)

2)A solid cylinder of radius 2.36 cm, length 4.98 cm, emissivity 0.839, and temperature 33.9°C is suspended in an environment of temperature 50.7°C.

a. Calculate the cylinder’s net thermal radiation transfer rate( Pnet 1)

b. If the cylinder is stretched until its radius is 0.426 cm, its net thermal radiation transfer rate becomes Pnet 2. Determine the ratio Pnet 2/Pnet

 



1
Expert's answer
2021-06-16T15:38:24-0400

1)

"(c_c M+c_wm) (t_2-t_1)=L\Delta m+c_w\Delta m(t_3-t_2),\implies"

"L=\frac{(c_cM+c_wm)(t_2-t_1)-c_w\Delta m(t_3-t_2)}{\Delta m}\approx 2.7\frac{MJ}{kg}."

2)

"A_1=2\pi r_1^2+2\pi r_1h_1=0.01088~m^2,"

"P_1=\sigma \epsilon (T_2^4-T_1^4)=1.0926\frac{W}{m^2},"

"V_2=V_1,\implies h_2=h_1\frac{r_1^2}{r_2^2}=152,84~cm,"

"A_2=2\pi r_2^2+2\pi r_2h_2=0.041~m^2,"

"\frac{P_2}{P_1}=\frac{A_2}{A_1}=3.77."